Question 15 of 15 incorrect ncorech incorrect inco
Answer #1

a) Al \rightarrow Al3+ + 3e-                 E= +0.67V

Ag+ + e- \rightarrow Ag                    E= +0.80V

Zn \rightarrow Zn2+ + 2e-                   E= +0.76V

Ni2+ + 2e-\rightarrow Ni                    E= -0.25V
......... ............... ........         E = +1.98V total potential

Ni and Ag plated out.

b) Ni and Ag are plated out; the potential is the same i.e E=+1.98V

c) Atomic weights: Ag=108, Ni=59

= 4.70gAg\; \ast \; \frac{1molAg}{108gAg}\ast \frac{2molNi}{1molAg}\; \ast \frac{59gNi}{1molNi} = 5.13g\: Ni